OpenStudy (anonymous):

y=1/4tan(6x-4) find second derivative

OpenStudy (anonymous):

is it \[y = \frac{ 1 }{ 4 }\tan(6x-4)\] ?????????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay we will use chain rule over here \[\frac{ dy }{ dx} = \sec^2 (6x-4)*\frac{ d }{ dx }(6x-4)\]

OpenStudy (anonymous):

this is the first derivative did u follow the step?

OpenStudy (anonymous):

yes thank you for helping me out

OpenStudy (anonymous):

in the previous step i forgot to put 1/4 \[\frac{ dy }{ dx } = \frac{ 1 }{ 4 }* \sec^2(6x-4) * 6\]

OpenStudy (anonymous):

now take the second derivative and show what u got?

OpenStudy (anonymous):

18sec^2(6x-4)tan(6x-4)

OpenStudy (anonymous):

yeah u r right

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

pleasure

OpenStudy (anonymous):

I have other question.:).y^6+x^3=y^2+10x (0,1)